# Wider or Deeper? Scaling LLM Inference-Time Compute with Adaptive Branching Tree Search

### v3

@@ Appendix D Multi-LLM AB-MCTS
Depending on the task, using more than one LLM for answer search can be advantageous. For example, if LLM A can generate more diverse initial answers and LLM B excels at refinement, building the answer tree with both models is expected to improve performance. This section demonstrates how AB-MCTS can be extended to scenarios in which multiple LLMs are available for answer generation, and also reports the experimental results for ARC-AGI-2, demonstrating the effectiveness of the proposed method.
@@ D.1.2 Generator Selection Algorithm I: Single GEN Node
Multi-LLM AB-MCTS-A with Generator Selection Algorithm I. In this algorithm, Multi-LLM AB-MCTS-A assigns each $l$-th generator an independent prior–Gaussian (Equations 11-16) or Beta (Equations 17-19)–depending on the score metric, and updates these priors to posteriors using the scores of the nodes in $\mathcal{N}_{l}$. After calculating the posteriors for all $l$, we perform Thompson sampling to choose a generator.
@@ D.2.1 Experimental Setup
In this experiment, we evaluated our approach on the 120 problems comprising the public evaluation set of ARC-AGI-2. For each problem, the generation budget was set to 250. The solution generation and refinement procedures are the same as those of ARC-AGI-1 experiment: The models were instructed to generate the transformation rule as Python code, and the search was guided by a reward signal corresponding to the number of demonstration cases correctly solved by the generated code. For solution generation, we used three frontier reasoning models: Gemini-2.5 Pro (gemini-2.5-pro-preview-05-06[44], o4-mini (o4-mini-2025-04-16[45], and DeepSeek-R1-0528 (deepseek-r1-0528[46]. We set the temperature to 0.6 for all models.

### v5

@@ Appendix D Multi-LLM AB-MCTS
Depending on the task, using more than one LLM for answer search [43] can be advantageous. For example, if LLM A can generate more diverse initial answers and LLM B excels at refinement, building the answer tree with both models is expected to improve performance. This section demonstrates how AB-MCTS can be extended to scenarios in which multiple LLMs are available for answer generation, and also reports the experimental results for ARC-AGI-2, demonstrating the effectiveness of the proposed method.
@@ D.1.2 Generator Selection Algorithm I: Single GEN Node
Multi-LLM AB-MCTS-A with Generator Selection Algorithm I. In this algorithm, Multi-LLM AB-MCTS-A assigns each $l$-th generator an independent prior–Gaussian (Equations 12-17) or Beta (Equations 18-20)–depending on the score metric, and updates these priors to posteriors using the scores of the nodes in $\mathcal{N}_{l}$. After calculating the posteriors for all $l$, we perform Thompson sampling to choose a generator.
@@ D.2.1 Experimental Setup
In this experiment, we evaluated our approach on the 120 problems comprising the public evaluation set of ARC-AGI-2. For each problem, the generation budget was set to 250. The solution generation and refinement procedures are the same as those of ARC-AGI-1 experiment: The models were instructed to generate the transformation rule as Python code, and the search was guided by a reward signal corresponding to the number of demonstration cases correctly solved by the generated code. For solution generation, we used three frontier reasoning models: Gemini-2.5 Pro (gemini-2.5-pro-preview-05-06[45], o4-mini (o4-mini-2025-04-16[46], and DeepSeek-R1-0528 (deepseek-r1-0528[22]. The temperature is set to be $0.6$ for all the models.